Problem 2.182, page 100. Do part (a). Ignore part (b) in the text, and continue with the following questions (b) Pick site A, what proportion of artifacts were less than 220 cm from the central hearth? c) What proportion of artifacts were greater that or equal to 60 cm from the central hearth? (d) What proportion of artifacts were in between 100 cm and 180 cm from the central hearth? e) What proportion of artifacts were less than 160 cm from the central hearth? (Careful !)8 Site 40 80 120 160 200 240 280 320 360 400 440 480 520 Distance (cm) I am stuck with the last question, that what's the proportion of less than 160 cm ?
Answer #1

e)
we need to know proportion of artifacts that were less than 160 cm ,
observe that 160 is the mid-point.

we will taken half of frequency of that bar { 6/2 }
P(X < 160) = 1 + 2 + 7 + 6/2
= 13
total number = 1 + 2 + 7 + 6 + 4 + 2 + 3 + 4*4 + 7 + 3 =51
required proportion = 13/51